OpenStudy (anonymous):

What value of n solves the equation? 3^n = 1/81 A. –243 B. –27 C. –4 D. 4

OpenStudy (mathstudent55):

Can you write 81 as a power of 3?

OpenStudy (anonymous):

I honestly don't know, mathstudent55.

OpenStudy (anonymous):

Could you tell me how?

OpenStudy (mathstudent55):

\(3^1 = 3\) \(3^2 = 3 \times 3 = 9\) \(3^3 = 3 \times 3 \times 3 = 27\) \(3^4 = 3 \times 3 \times 3 \times 3 = 81\) Do you see that \(81 = 3^4\) ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

I get it now.

OpenStudy (mathstudent55):

Now look at the right side of the original problem. You have \(\dfrac{1}{81} \) Since we now know that \(81 = 3^4\), we can rewrite the right siuda as \(\dfrac{1}{81} = \dfrac{1}{3^4} \)

OpenStudy (anonymous):

Okay.

OpenStudy (mathstudent55):

The question now is: \(3^n = \dfrac{1}{3^4} \) Now we use the definition of a negative exponent: \(a^{-n} = \dfrac{1}{a^n} \)

OpenStudy (mathstudent55):

We start from the right side of the definition: \(\dfrac{1}{a^n} = a^{-n} \) If we use the fraction from your problem, and we follow the definition, we get: \(\dfrac{1}{3^4} = 3^{-4} \)

OpenStudy (mathstudent55):

The exponent, n, is -4. Answer: C. -4

OpenStudy (anonymous):

I get it now. Thanks!